OpenStudy (anonymous):

Which of the following number lines represents the solution to m/-2 - 7 = -9? 11. -4. 4. 1.

OpenStudy (anonymous):

people answer it lol

OpenStudy (anonymous):

see m /-2 -7 = -9 so can be arranged as -m/2 -7 = -9 -m/2 = -9 +7 -m/2 = -2 -m = -4 m =4 be ans

OpenStudy (anonymous):

look at the smart a$$ over here lol

OpenStudy (anonymous):

THANKS SO MUCH!!

OpenStudy (anonymous):

put 15 up